Required information

[The following information applies to the questions displayed below.]

Adger Corporation is a service company that measures its output based on the number of customers served. The company provided the following fixed and variable cost estimates that it uses for budgeting purposes and the actual results for May as shown below:

Fixed Element
per Month
Variable Element per Customer Served Actual Total
for May
Revenue $ 5,600 $ 182,000
Employee salaries and wages $ 55,000 $ 1,600 $ 110,300
Travel expenses $ 850 $ 27,200
Other expenses $ 34,000 $ 32,600

When preparing its planning budget the company estimated that it would serve 30 customers per month; however, during May the company actually served 35 customers.

Required:

1. What amount of revenue would be included in Adger’s flexible budget for May?

2. What amount of employee salaries and wages would be included in Adger’s flexible budget for May?

3. What amount of travel expenses would be included in Adger’s flexible budget for May?

4. What amount of other expenses would be included in Adger’s flexible budget for May?

5. What net operating income would appear in Adger’s flexible budget for May?

6. What is Adger’s revenue variance for May? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

7. What is Adger’s employee salaries and wages spending variance for May? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

8. What is Adger’s travel expenses spending variance for May? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

9. What is Adger’s other expenses spending variance for May? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

10. What amount of revenue would be included in Adger’s planning budget for May?

11. What amount of employee salaries and wages would be included in Adger’s planning budget for May?

12. What amount of travel expenses would be included in Adger’s planning budget for May?

13. What amount of other expenses would be included in Adger’s planning budget for May?

14. What activity variance would Adger report in May with respect to its revenue? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)

15. What activity variances would Adger report with respect to each of its expenses for May? (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1

1) Revenue in Flexible budget = 35*5600 = 196000

2)Employee salaries and wages in flexible budget = (35*1600+55000) = 111000

3) Travel expense in Flexible budget = 35*850 = 29750

4) Other expense in Flexible budget = 34000

Note : Please post whole question in the group of 4-4 sub parts as per HOMEWORKLIB RULES